OpenStudy (natalie1234):

Will give medal If x varies inversely with y and x = 8 when y = 6, find y when x = 10

OpenStudy (anonymous):

do you remember the formula for inverse variation?

OpenStudy (anonymous):

y=k/x

OpenStudy (natalie1234):

umm y=k/x?

OpenStudy (anonymous):

so... 6=k/8.... so solve for k now

OpenStudy (anonymous):

k=48... so now, y=k/x... would be y=48/10, which gives you 4.8=y

OpenStudy (natalie1234):

ohh ok thanks (:

OpenStudy (anonymous):

your welcome :) thank u :)